### v2

# Sequential Randomization Tests Using E-values: A Betting Approach for Clinical Trials

Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We present a nonparametric sequential test – the randomization e-process (e-RT) – that derives validity solely from the randomization mechanism. Using a betting framework, e-RT constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. The e-RT provides a conservative, assumption-free complement to model-based sequential analyses.
Keywords: e-values, e-process, randomization test, sequential analysis, clinical trials
1 Introduction
Sequential monitoring of randomized controlled trials requires methods that control Type I errors regardless of when or why the monitoring stops. Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries. When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guarantees may erode.
E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025). An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1. This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
We present the randomization e-process (e-RT), a sequential test that derives validity solely from the trial’s randomization mechanism. The construction uses a betting framework: at each enrollment, we wager on the treatment assignment given the observed outcome. Under the null hypothesis of no treatment effect, treatment assignment is independent of outcome, and no betting strategy can systematically grow wealth. This independence, guaranteed by randomization, is the only assumption required.
The contributions of this note are as follows.
  1. 1.
    An explicit construction of e-RT with proof of validity
  2. 2.
    Simulation studies demonstrating Type I error calibration and power
2 The e-RT Construction
2.1 Setup
Consider a sequential randomized trial. At each enrollment $i=1,2,\ldots$, we observe $T_{i}$: assignment of treatment (control or intervention) and $Y_{i}$: binary outcome (event or no event). Treatment is assigned with known probability $p$, typically $p=0.5$ for 1:1 randomization.
The null hypothesis is that treatment has no effect. Under this hypothesis, knowing the outcome of a patient tells you nothing about which arm they were assigned to – outcome and treatment are independent ($Y_{i}\perp T_{i}$).
2.2 Betting Construction
e-RT constructs a wealth process $W_{n}$ by betting sequentially on treatment assignments.
For each patient, after observing their outcome but before learning their treatment assignment, we place a bet. The bet is a fraction $\lambda_{i}\in[0,1]$ representing how much we wager on the patient being in the treatment arm.
Under the null hypothesis, outcome tells us nothing about assignment; hence we can’t systematically win. Under the alternative, outcomes differ by arm, so we can bet profitably.
The wealth updates as:
$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=\text{intervention}\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=\text{control}\end{cases}$ (1)
starting from $W_{0}=1$. Note that the intervention and control here are simply labels, they could be called whatever (A versus B, etc).
2.3 Adaptive Betting Strategy
The betting strategy uses accumulated data to inform each bet. Here is how it works:
  1. 1.
    A patient is randomized (assignment blinded).
  2. 2.
    Outcome is observed.
  3. 3.
    Historical records are checked. The most likely arm for this outcome is defined.
  4. 4.
    Bet is placed according to previous data.
  5. 5.
    Assignment is revealed.
  6. 6.
    Wealth updated. If guess is correct, it grows; if wrong, it shrinks.
Under the null hypothesis, any historical imbalance between arms is noise. Bets will be sometimes right, sometimes wrong, and wealth fluctuates randomly around 1. Under the alternative, the imbalance is real signal. Bets will be systematically correct, and wealth grows over time.
Formally, let:
$\hat{\delta}_{i-1}=\text{(event rate in treatment)}-\text{(event rate in control)}$ (2)
estimated from all patients before $i$. The betting fraction is:
$\lambda_{i}=\begin{cases}0.5+0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{event}\\ 0.5-0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{no event}\end{cases}$ (3)
where $c_{i}\in[0,1]$ is a confidence parameter that ramps up from 0 to 1 over a burn-in period, preventing wild bets when data are sparse:
$c_{i}=\min\left(1,\frac{i-\text{burn-in}}{\text{ramp}}\right)$ (4)
The above can be read as: At any specific point, the betting fraction $\lambda$ is defined as $0.5$ plus the wager if there has been an event and $0.5$ minus the wager if there hasn’t been an event. The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{i}$. This factor is simply a factor that avoids us from making full bets early on. It is determined by a burn-in period (where we don’t bet, just collect data), and a ramp, which is also arbitrarily defined, so that wager is "muffled" early on. Finally, the $0.5$ multiplier before $c_{i}$ and the wager ($\hat{\delta}_{i-1}$) simply keeps $\lambda$ between zero and one.
2.4 Worked Example
Consider a trial comparing intervention versus control, with a binary outcome (event or no event). Event is mortality, which is expected to be lower with intervention. Allocation is 1:1 ($p=0.5$). Assume burn-in is complete ($c=1$).
We look back at patients 1–199. Intervention arm has 100 patients, 35 events, so rate = 35.0%. Control arm has 99 patients, 40 events, so rate = 40.4%. $\hat{\delta}_{199}=0.350-0.404=-0.054$ (intervention looks protective).
Patient 200 has an event (dies). Where is this patient likely from? Events are more common in control (40.4% vs 35.0%), so probably control. We bet $\lambda=0.5+0.5\times(-0.054)=0.473$ on intervention and $1-\lambda=0.527$ on control. Assignment revealed and patient indeed is in control group. We guessed right! Our multiplier will be $0.527/0.5=1.054$, hence wealth grows 5.4%.
Now patient 201 has been enrolled. First, we update the counts with patient 200’s data. Intervention arm now has 100 patients, 35 events (rate = 35.0%), and control arm has 100 patients, 41 events (rate = 41.0%). The $\hat{\delta}_{200}=0.350-0.410=-0.060$. Patient 201 does not have an event. Non-events are more common in intervention (65.0% vs 59.0%), so probably intervention. We will bet once again. This time we bet $\lambda=0.5-0.5\times(-0.060)=0.530$ on intervention and $1-\lambda=0.470$ on control. Finally, assignment is revealed and we guessed right again: patient is in intervention. Our multiplier =0.530/0.5=1.060=0.530/0.5=1.060. Wealth grows another 6.0%.
Finally, patient 202 has been enrolled. We update counts once more with intervention having 101 patients and 35 events (rate = 34.7%) and control having 100 patients with 41 events (rate = 41.0%). Our $\hat{\delta}_{201}=0.347-0.410=-0.063$. Patient 202 has an event (dies). Events cluster in control, so we bet toward control: $\lambda=0.5+0.5\times(-0.063)=0.469$ on intervention and $1-\lambda=0.531$ on control. Assignment is revealed and we guessed wrong—patient is in intervention group! Our multiplier will be $0.469/0.5=0.938$. Wealth shrinks 6.2%.
Cumulative wealth is the product of all multipliers:
$W_{202}=W_{199}\times 1.054\times 1.060\times 0.938=W_{199}\times 1.048$ (5)
Despite one wrong guess, wealth grew 4.8% over these three patients.
Even when intervention works, some intervention patients die. We guess wrong sometimes. But if intervention truly reduces mortality, events cluster in control and non-events cluster in intervention — we are right more often than wrong, and wealth grows on average.
Under the null hypothesis (no effect), right and wrong guesses balance out. Wealth fluctuates randomly around 1.
In summary: the outcome determines $\lambda$ (which direction to bet); the treatment arm determines which multiplier formula to apply (whether the bet paid off). The machinery is agnostic to whether the effect size is following the expected direction; it simply keeps betting according to what is more likely.
2.5 Validity
Theorem 1.
Under the null hypothesis, the wealth process $(W_{n})$ is a test martingale — its expected value cannot grow.
Proof.
Under the null, outcome and treatment are independent. So after seeing the outcome, treatment assignment is still just a coin flip with probability $p$. The expected wealth multiplier is:
$\mathbb{E}[\text{multiplier}]$ $=p\times\frac{\lambda}{p}+(1-p)\times\frac{1-\lambda}{1-p}$ (6)
$=\lambda+(1-\lambda)$ (7)
$=1$ (8)
No matter how we choose our bet $\lambda$, we can’t systematically grow wealth when the null is true. ∎
Corollary 1.
For any stopping time $\tau$, $\mathbb{E}[W_{\tau}]\leq 1$ under the null. By Markov’s inequality, $P(W_{\tau}\geq 1/\alpha)\leq\alpha$. Thus rejecting when $W_{\tau}\geq 1/\alpha$ controls Type I error at level $\alpha$, regardless of the stopping rule.
3 Simulation Studies
We evaluated e-RT operating characteristics by simulation. For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size. We used burn-in = 50 patients and ramp = 100 patients. Control arm event rate was 40% in all scenarios.
3.1 Results
Table 1 presents Type I error and power for trials designed to detect 5% or 10% absolute risk reductions (ARR) with 80% or 90% power.
Table 1: e-RT Operating Characteristics
ARR Target Power $n$ Type I Error e-RT Power Median Crossing
5% 80% 2942 0.032 48.6% 1392 (47%)
10% 80% 712 0.021 50.4% 401 (56%)
5% 90% 3938 0.035 62.8% 1842 (47%)
10% 90% 954 0.025 65.9% 478 (50%)
Type I error was well controlled across all scenarios (0.021 – 0.035), below the nominal $\alpha=0.05$. This confirms the theoretical guarantee from the martingale property.
Power was approximately 50% for trials designed with 80% power, and 63–66% for trials designed with 90% power. When the e-RT rejected the null, it did so at approximately half the planned sample size (median crossing 47–56% of total enrollment).
3.2 Interpretation
The e-RT is not a replacement for the primary analysis—it is a continuous monitoring tool. A trial designed for 80% power with a traditional test gains a 50% chance of stopping early via e-RT, at roughly the halfway point. If the threshold is not crossed, the trial proceeds to completion and the planned analysis is conducted.
This represents a free option: anytime-valid monitoring with no alpha spending and no pre-specified interim looks. The “cost” is that the e-RT alone has lower power than a fixed-sample test. But when layered on top of a properly powered trial, it provides early stopping when effects are larger than anticipated.
3.3 Trajectory Examples
Figure 1 shows example e-RT trajectories from 30 simulated trials under the null hypothesis (both arms 40% event rate). Sample sizes correspond to trials designed for 80% power (n = 712, left) and 90% power (n = 954, right) to detect a 10% ARR. Under the null, wealth fluctuates randomly around 1. Some trajectories temporarily rise toward the threshold but eventually drift back down. The downward drift over time reflects the accumulating “cost” of betting on noise.
Figure 1: e-RT trajectories under the null hypothesis. Left: n = 712 (80% power design). Right: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Dotted gray line: neutral (wealth = 1). Under the null, no trajectory crosses the threshold.
Figure 2 shows trajectories under the alternative hypothesis (40% vs 30% event rates, true ARR = 10%). With a real treatment effect, wealth grows systematically. Most trajectories cross the rejection threshold before enrollment completes, and many reach values far exceeding 20 — providing strong evidence against the null.
Figure 2: e-RT trajectories under the alternative hypothesis (true ARR = 10%). Panel A: n = 712 (80% power design). Panel B: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Under the alternative, approximately half of trajectories cross the threshold, typically around the midpoint of enrollment.
4 Discussion
The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values. The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses.
4.1 Operating characteristics
Simulations demonstrate that e-RT controls Type I error conservatively (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
These results should be interpreted carefully. The e-RT is not designed to replace the primary analysis of a trial; at least not yet. Rather, it provides a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule. If the e-RT crosses its threshold, one could consider stopping the trial. If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
4.2 Relationship to existing work
The betting framework for hypothesis testing was developed by shafer2021. E-values and e-processes have been extensively studied (vovk2021; ramdas2022; ramdas2025). koning2025 develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations. Our construction differs in using patient-by-patient betting with adaptive learning of the effect size. These approaches share theoretical foundations but yield different procedures.
4.3 Limitations
Several limitations should be noted. This is an experimental method under development. So far, this has only been tested for binary outcomes. It is reasonable to assume that extending to continuous endpoints would be feasible; this is planned. It is uncertain how this method could be applied for time-to-event endpoints. Hypothetically, we could hold the bet value for all patients randomized in a period until outcomes arrive, and approaches such as restricted mean survival time could be used. This is also under development.
There are other limitations. The e-RT tests only whether event rates differ between arms. It does not directly estimate treatment effects or provide confidence intervals. The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates. Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power. Fourth, our simulations used a specific betting strategy (burn-in = 50, ramp = 100); other choices may yield different operating characteristics. Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power. Finally, it is unclear how this method will behave in situations where heterogeneity in treatment effects exist or there are temporal instabilities in effect size. The method could be extended to bet according to relative effect size approaches, such as the odds ratio. This is also under development.
4.4 Conclusion
The e-RT provides anytime-valid sequential inference for randomized trials using only the guarantee of randomization. Its validity is unconditional on the data-generating process, making it a robust tool for trial monitoring. While it trades power for this robustness, it offers a valuable complement to traditional analysis methods.
4.5 Disclaimer
This is an experimental method under development. Application to real patients should only be considered under surveillance from an experienced statistician. The author is not responsible for consequences of use of this method.
4.6 Version Control
  1. 1.
    First Version (Dec 04, 2025)
  2. 2.
    Second Version (this version): Minor text adjustments; removed claim on sharp null.
Appendix A R Code
# e-RT Simulations: Type I Error and Power (Unified)
library(tidyverse)
# Core e-RT function
compute_eRT <- function(treatment, outcome, p = 0.5, burn_in = 50, ramp = 100) {
n <- length(treatment)
if (is.factor(treatment)) treatment <- as.numeric(treatment) - 1
if (is.factor(outcome)) outcome <- as.numeric(outcome) - 1
wealth <- numeric(n)
wealth[1] <- 1
for (i in 2:n) {
trt_prev <- treatment[1:(i-1)]
out_prev <- outcome[1:(i-1)]
rate_trt <- mean(out_prev[trt_prev == 1])
rate_ctrl <- mean(out_prev[trt_prev == 0])
if (is.nan(rate_trt)) rate_trt <- 0.5
if (is.nan(rate_ctrl)) rate_ctrl <- 0.5
delta_hat <- rate_trt - rate_ctrl
c_i <- max(0, min(1, (i - burn_in) / ramp))
if (outcome[i] == 1) {
lambda <- 0.5 + 0.5 * c_i * delta_hat
} else {
lambda <- 0.5 - 0.5 * c_i * delta_hat
lambda <- max(0.001, min(0.999, lambda))
if (treatment[i] == 1) {
multiplier <- lambda / p
} else {
multiplier <- (1 - lambda) / (1 - p)
wealth[i] <- wealth[i-1] * multiplier
return(wealth)
# Simulate a single trial
simulate_trial <- function(n, p_trt = 0.5, rate_trt, rate_ctrl) {
treatment <- rbinom(n, 1, p_trt)
outcome <- numeric(n)
outcome[treatment == 1] <- rbinom(sum(treatment == 1), 1, rate_trt)
outcome[treatment == 0] <- rbinom(sum(treatment == 0), 1, rate_ctrl)
return(data.frame(treatment = treatment, outcome = outcome))
# Unified simulation function
simulate_eRT <- function(n_sims = 5000,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100) {
# Determine hypothesized ARR for sample size calculation
if (is.null(hypothesized_ARR) && is.null(p_trt)) {
stop("Mustspecifyeitherp_trtorhypothesized_ARR")
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
# Calculate sample size based on hypothesized effect
$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
power = target_power, sig.level = alpha)
n_per_arm <- ceiling(ss\$n)
n_patients <- 2 * n_per_arm
# Determine true p_trt (null if not specified)
if (is.null(p_trt)) {
p_trt <- p_ctrl # Null is true
true_ARR <- 0
sim_type <- "TypeIError"
} else {
true_ARR <- p_ctrl - p_trt
sim_type <- "Power"
[…]
p_ctrl, hypothesized_ARR, target_power * 100))
[…]
p_ctrl, p_trt, true_ARR))
[…]
rejections <- 0
first_crossing <- numeric(n_sims)
final_evalues <- numeric(n_sims)
pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
for (sim in 1:n_sims) {
trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
final_evalues[sim] <- wealth[n_patients]
crossing <- which(wealth >= 1/alpha)
if (length(crossing) > 0) {
rejections <- rejections + 1
first_crossing[sim] <- crossing[1]
} else {
first_crossing[sim] <- NA
setTxtProgressBar(pb, sim)
close(pb)
rejection_rate <- rejections / n_sims
se <- sqrt(rejection_rate * (1 - rejection_rate) / n_sims)
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
if (sim_type == "Power") {
[…]
} else {
[…]
return(list(
sim_type = sim_type,
rejection_rate = rejection_rate,
se = se,
n_per_arm = n_per_arm,
n_patients = n_patients,
p_ctrl = p_ctrl,
p_trt = p_trt,
hypothesized_ARR = hypothesized_ARR,
true_ARR = true_ARR,
target_power = target_power,
alpha = alpha,
median_crossing = median_stop,
first_crossing = first_crossing,
final_evalues = final_evalues
# Plot trajectories
plot_trajectories <- function(n_trials = 30,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100,
title = NULL) {
if (is.null(hypothesized_ARR) && is.null(p_trt)) {
stop("Mustspecifyeitherp_trtorhypothesized_ARR")
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
if (is.null(p_trt)) {
p_trt <- p_ctrl # Null
# Calculate sample size
$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
power = target_power, sig.level = alpha)
n_patients <- 2 * ceiling(ss\$n)
true_ARR <- p_ctrl - p_trt
if (is.null(title)) {
if (true_ARR == 0) {
title <- sprintf("e-RTUnderNull(n=%d,designedfor%.0f%%ARR)",
n_patients, hypothesized_ARR * 100)
} else {
title <- sprintf("e-RTUnderAlternative(n=%d,%.0f%%power)",
n_patients, target_power * 100)
trajectories <- list()
for (i in 1:n_trials) {
trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
trajectories[[i]] <- data.frame(
patient = 1:n_patients,
wealth = wealth,
trial = i
df <- bind_rows(trajectories)
p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
geom_line(alpha = 0.4, color = "steelblue") +
geom_hline(yintercept = 1/alpha, linetype = "dashed", color = "red") +
geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
scale_y_log10() +
labs(
title = title,
subtitle = sprintf("Control=%.0f%%,Treatment=%.0f%%,TrueARR=%.0f%%",
p_ctrl * 100, p_trt * 100, true_ARR * 100),
x = "Patient",
y = "e-value(logscale)"
annotate("text", x = n_patients * 0.95, y = 1/alpha * 1.5,
label = sprintf("1/alpha=%.0f", 1/alpha), color = "red", hjust = 1) +
theme_minimal() +
theme(
plot.title = element_text(face = "bold"),
panel.grid.minor = element_blank()
return(p)
# Run all simulations
run_all <- function(n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05,
seed = 42) {
set.seed(seed)
results <- list()
# Create grid of scenarios
scenarios <- expand.grid(
hypothesized_ARR = ARRs,
target_power = target_powers,
stringsAsFactors = FALSE
# For each scenario, run Type I error and Power
all_results <- data.frame()
for (i in 1:nrow(scenarios)) {
hyp_ARR <- scenarios\$hypothesized_ARR[i]
tgt_pow <- scenarios\$target_power[i]
[…]
hyp_ARR * 100, tgt_pow * 100))
[…]
# Type I error (null true, same sample size)
[…]
t1 <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = NULL, # Null is true
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
alpha = alpha
# Power (alternative true)
[…]
pow <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = p_ctrl - hyp_ARR, # True effect = hypothesized
target_power = tgt_pow,
alpha = alpha
all_results <- rbind(all_results, data.frame(
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
n_patients = t1\$n_patients,
type1_error = t1\$rejection_rate,
type1_se = t1\$se,
eRT_power = pow\$rejection_rate,
power_se = pow\$se,
median_crossing = pow\$median_crossing
results\$summary <- all_results
# Print summary
[…]
print(all_results %>%
mutate(
hypothesized_ARR = sprintf("%.0f%%", hypothesized_ARR * 100),
target_power = sprintf("%.0f%%", target_power * 100),
type1_error = sprintf("%.3f", type1_error),
eRT_power = sprintf("%.1f%%", eRT_power * 100)
select(hypothesized_ARR, target_power, n_patients,
type1_error, eRT_power, median_crossing))
return(results)
# Run
if (interactive()) {
results <- run_all(
n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05
# Trajectory plots - Alternative
p1 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.80)
ggsave("traj_alt_10pct_80pow.pdf", p1, width = 8, height = 5)
p2 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.90)
ggsave("traj_alt_10pct_90pow.pdf", p2, width = 8, height = 5)
# Trajectory plots - Null
p3 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.80)
ggsave("traj_null_10pct_80pow.pdf", p3, width = 8, height = 5)
p4 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.90)
ggsave("traj_null_10pct_90pow.pdf", p4, width = 8, height = 5)