### v4

# Sequential Randomization Tests Using e-values: Applications for trial monitoring

Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We discuss a nonparametric sequential test and its application to continuous and time-to-event endpoints that derives validity solely from the randomization mechanism. Using a betting framework, these tests constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. These methods provide a conservative, assumption-free complement to model-based sequential analyses.
Keywords: e-values, e-process, randomization test, sequential analysis, clinical trials
1 Introduction
Sequential monitoring of randomized controlled trials requires methods that control Type I errors regardless of when or why the monitoring stops. Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries. When these assumptions fail, or when trials adapt in ways not fully prespecified, validity guarantees may erode.
Monitoring clinical trials, specially in the context of acutely ill patients, is of paramount importance. Traditional methods include $\alpha-spending$ functions. When interim analyses are spaced, evidence grows unnoticed between the interim analyses. This may both delay the implementation of potentially useful strategies or prolong trials when safety signals arise early.
E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025). An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1. This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which tests treatment effects by wagering on treatment assignments given observed outcomes. The intuition is discussed by ramdas2021game. Under the null hypothesis, randomization ensures that assignments are independent of outcomes, so no betting strategy can systematically accumulate wealth.
We propose e-RT (e-value Randomized Trial), a method for prospective sequential monitoring of randomized trials. Like i-bet (pmlr-v177-duan22a), e-RT uses betting martingales for inference, but differs in key respects: e-RT monitors sequentially as patients enroll rather than analyzing completed trial data; it requires no covariates or working models; and it uses an adaptive betting fraction derived from running outcome estimates rather than fixed wagers. This yields a method with minimal assumptions suitable for real-time trial monitoring.
2 Overall construction
2.1 Setup
Consider a sequential randomized trial with 1:1 allocation. At each enrollment $i=1,2,\ldots$, we observe:
  • $T_{i}\in\{0,1\}$: treatment assignment (0 = control, 1 = intervention)
  • $Y_{i}\in\{0,1\}$: binary outcome (0 = no event, 1 = event)
Treatment is assigned with known probability $p=P(T_{i}=1)$, typically $p=0.5$.
The null hypothesis is that treatment assignment has no effect on outcome:
$H_{0}:Y_{i}\perp T_{i}\text{ for all }i$ (1)
Under this hypothesis, observing the outcome provides no information about which arm the patient was assigned to.
2.2 Wealth Process
Following pmlr-v177-duan22a, we construct a wealth process by wagering on treatment assignments. After observing outcome $Y_{i}$ but before learning treatment assignment $T_{i}$, we choose $\lambda_{i}\in[0,1]$: the fraction wagered on intervention.
The wealth updates as:
$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=1\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=0\end{cases}$ (2)
starting from $W_{0}=1$. When we bet toward the correct arm, wealth grows; when wrong, it shrinks.
2.3 Betting Strategy
The validity guarantee holds for any betting strategy where $\lambda_{i}$ depends only on $\mathcal{F}_{i-1}$. Power depends on choosing bets that grow wealth under the alternative. We use a strategy that learns the treatment effect from accumulating data.
Let:
$\hat{\delta}_{i-1}=\text{(event rate in intervention)}-\text{(event rate in control)}$ (3)
estimated from patients $1,\ldots,i-1$. The betting fraction is:
$\lambda_{i}=\begin{cases}0.5+0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=1\\ 0.5-0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=0\end{cases}$ (4)
where $c_{i}\in[0,1]$ ramps from 0 to 1 over a burn-in period:
$c_{i}=\min\left(1,\max\left(0,\frac{i-n_{0}}{n_{r}}\right)\right)$ (5)
with $n_{0}$ the burn-in period and $n_{r}$ the ramp period. This prevents large bets when $\hat{\delta}_{i-1}$ is unstable due to small samples.
The logic: if $\hat{\delta}>0$ (more events in intervention), then events suggest intervention and non-events suggest control. If $\hat{\delta}<0$ (fewer events in intervention), then events suggest control and non-events suggest intervention. The factor of $0.5$ before $c_{i}\cdot\hat{\delta}_{i-1}$ ensures $\lambda_{i}\in[0,1]$.
2.4 Worked Example
Consider a trial comparing intervention versus control, with a binary outcome (event or no event). Event is mortality, which is expected to be lower with intervention. Allocation is 1:1 ($p=0.5$). Assume burn-in is complete ($c_{i}=1$).
We look back at patients 1–199. Intervention arm has 100 patients, 35 events, so rate = 35.0%. Control arm has 99 patients, 40 events, so rate = 40.4%. $\hat{\delta}_{199}=0.350-0.404=-0.054$ (intervention looks protective).
Patient 200 has an event (dies). Where is this patient likely from? Events are more common in control (40.4% vs 35.0%), so probably control. We bet $\lambda=0.5+0.5\times(-0.054)=0.473$ on intervention and $1-\lambda=0.527$ on control. Assignment revealed: control. We guessed right. Multiplier: $0.527/0.5=1.054$. Wealth grows 5.4%.
Patient 201: Update counts—intervention has 100 patients, 35 events (35.0%); control has 100 patients, 41 events (41.0%). $\hat{\delta}_{200}=-0.060$. Patient 201 has no event. Non-events are more common in intervention (65.0% vs 59.0%), so probably intervention. Bet: $\lambda=0.5-0.5\times(-0.060)=0.530$. Assignment revealed: intervention. Multiplier: $0.530/0.5=1.060$. Wealth grows 6.0%.
Patient 202: Update counts—intervention has 101 patients, 35 events (34.7%); control has 100 patients, 41 events (41.0%). $\hat{\delta}_{201}=-0.063$. Patient 202 has an event. Bet toward control: $\lambda=0.5+0.5\times(-0.063)=0.469$. Assignment revealed: intervention. Wrong guess. Multiplier: $0.469/0.5=0.938$. Wealth shrinks 6.2%.
Cumulative wealth:
$W_{202}=W_{199}\times 1.054\times 1.060\times 0.938=W_{199}\times 1.048$ (6)
Despite one wrong guess, wealth grew 4.8% over these three patients. Under the alternative, correct guesses outnumber incorrect ones on average and wealth grows. Under the null, right and wrong guesses balance out and wealth fluctuates around 1.
2.5 Validity
Theorem 1.
Under the null hypothesis, the wealth process $(W_{n})$ is a nonnegative martingale.
Proof.
Under the null, outcome and treatment are independent. After observing outcome $Y_{i}$, treatment assignment remains a coin flip with $P(T_{i}=1)=p$. The expected multiplier given any bet $\lambda_{i}$ is:
$\mathbb{E}[\text{multiplier}\mid\lambda_{i}]$ $=p\times\frac{\lambda_{i}}{p}+(1-p)\times\frac{1-\lambda_{i}}{1-p}$ (7)
$=\lambda_{i}+(1-\lambda_{i})=1$ (8)
Thus $\mathbb{E}[W_{i}\mid W_{i-1},\lambda_{i}]=W_{i-1}$. ∎
Corollary 1.
By Ville’s inequality (ville1939), for any nonnegative martingale starting at 1:
$\Pr_{H_{0}}\left(\sup_{n\geq 1}W_{n}\geq\frac{1}{\alpha}\right)\leq\alpha$ (9)
Thus rejecting when wealth ever exceeds $1/\alpha$ controls Type I error at level $\alpha$, regardless of when or why monitoring stops.
3 Simulation Studies
We evaluated operating characteristics or e-RT by simulation. For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size. We used burn-in = 50 patients and ramp = 100 patients. Control arm event rate was 40% in all scenarios.
3.1 Results
Table 1 presents Type I error and power for trials designed to detect 5% or 10% absolute risk reductions (ARR) with 80% or 90% power.
Table 1: Operating Characteristics for Binary e-process
ARR Target Power $n$ Type I Error Binary e-process Power Median Crossing
5% 80% 2942 0.032 48.6% 1392 (47%)
10% 80% 712 0.021 50.4% 401 (56%)
5% 90% 3938 0.035 62.8% 1842 (47%)
10% 90% 954 0.025 65.9% 478 (50%)
Type I error was well controlled across all scenarios (0.021 – 0.035), below the nominal $\alpha=0.05$. This confirms the theoretical guarantee from the martingale property.
Power was approximately 50% for trials designed with 80% power, and 63–66% for trials designed with 90% power. When the process rejected the null, it did so at approximately half the planned sample size (median crossing 47–56% of total enrollment).
3.2 Interpretation
This method seems to be appealing as a continuous monitoring tool. A trial designed for 80% power with a traditional test gains a 50% chance of stopping early, at roughly the halfway point. If the threshold is not crossed, the trial proceeds to completion and the planned analysis is conducted with no penalty.
This represents a free option: anytime-valid monitoring with no alpha spending and no pre-specified interim looks. The “cost” is that this method alone has lower power than a fixed-sample test. But when layered on top of a properly powered trial, it provides early stopping when effects are larger than anticipated.
3.3 Trajectory Examples
Figure 1 shows example wealth trajectories from 30 simulated trials under the null hypothesis (both arms 40% event rate). Sample sizes correspond to trials designed for 80% power (n = 712, left) and 90% power (n = 954, right) to detect a 10% ARR. Under the null, wealth fluctuates randomly around 1. Some trajectories temporarily rise toward the threshold but eventually drift back down. The downward drift over time reflects the accumulating “cost” of betting on noise.
Figure 1: Wealth trajectories under the null hypothesis. Left: n = 712 (80% power design). Right: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Dotted gray line: neutral (wealth = 1). Under the null, no trajectory crosses the threshold.
Figure 2 shows trajectories under the alternative hypothesis (40% vs 30% event rates, true ARR = 10%). With a real treatment effect, wealth grows systematically. Most trajectories cross the rejection threshold before enrollment completes, and many reach values far exceeding 20 — providing strong evidence against the null.
Figure 2: E-processes trajectories under the alternative hypothesis (true ARR = 10%). Left: n = 712 (80% power design). Right: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Under the alternative, approximately half of trajectories cross the threshold, typically around the midpoint of enrollment.
4 Continuous Outcomes
The e-RT treats each patient as a single Bernoulli trial: the outcome (event vs. no event) is observed, and we bet on which arm that patient came from. For continuous endpoints, the logic is the same but the signal is richer. Each patient now contributes a continuous measurement (for example, ventilator-free days, change in biomarker, or a physiologic score), and the betting strategy uses how extreme that value is relative to past data. Therefore, defining wager is slightly more granular.
One may extend this to continuous endpoint which we shall call e-RTC. The validity still comes only from randomization: under the null hypothesis, the distribution of the continuous outcome is the same in both arms, so the outcome does not help to predict treatment assignment.
4.1 Setup
At each enrollment $i=1,2,\ldots$, the trial generates two pieces of information. First, the randomization mechanism assigns the patient to one of the two arms, denoted by $T_{i}\in\{0,1\}$ (with 0 for control and $1$ for intervention), using a known allocation probability $p$ (typically $p=0.5$ for 1:1 randomization, although equal allocation is not required). Second, once follow-up is complete, we observe a continuous outcome $Y_{i}\in\mathbb{R}$ such as ventilator–free days, a biomarker concentration, or a physiologic measurement. So far, same as before but using a continuous endpoint.
Under the null hypothesis of no treatment effect, the distribution of $Y_{i}$ is identical in both arms, and hence the outcome carries no information about the treatment assignment; formally, $Y_{i}\perp T_{i}$ under $H_{0}$. This single independence relationship is the foundation of the continuous randomization e-process. The idea mirrors the binary approach: each patient creates a small betting game in which we observe $Y_{i}$, form a data-driven guess about which arm the patient is more likely to belong to, and then update our wealth once the actual assignment $T_{i}$ is revealed. If the null is true, these guesses cannot systematically win because outcomes are uninformative about treatment. If the alternative is true, outcomes begin to separate between arms, the bets gain predictive power, and wealth grows accordingly.
Viewed this way, extending the randomization e-process from binary to continuous outcomes may not require additional assumptions; it merely replaces the event/no-event signal with a continuous measure of extremeness relative to past observations, while preserving the anytime-valid martingale structure. The difference in how the wager is defined.
4.2 Betting strategy for continuous outcomes
We can use a data-driven rule with three ingredients:
  1. 1.
    A center and scale for the outcomes observed so far. Outcomes con be all all over the place. We need to standardize.
  2. 2.
    A standardized residual for the new outcome. As the reader will see we bet proportionally to how far the result is away from the center reference we chose. This also needs to be standardized.
  3. 3.
    A smooth map from that residual into a betting fraction $\lambda_{i}\in(0,1)$. If the measurement is an outlier, a huge wager would be placed because difference between measurement and the center of the scale would be huge, we need to muffle it to something between $(0,1)$.
In simpler terms: Each new outcome is first compared to the past outcomes to understand how “unusual” it is. To do this, we anchor the outcome to a robust center and a robust scale: the median gives the center and the MAD (median absolute deviation) gives the spread. We then compute a standardized value: how far above or below the median this new observation is, measured in MAD units. That standardized number is what guides how aggressively we bet. A large positive value means “this looks unusually high compared with past outcomes,” a large negative value means the opposite, and values near zero mean “this one looks typical.” The point of the MAD is simple: it behaves well even when early data are messy or skewed. It stops a single extreme value from blowing up the bet and keeps the e-process stable while the trial is still young. This likely comes at the cost of reduced power.
At step $i$, using all previous outcomes $Y_{1},\dots,Y_{i-1}$, we compute:
$m_{i-1}$ $=\text{median}(Y_{1},\dots,Y_{i-1}),$ (10)
$s_{i-1}$ $=\text{MAD}(Y_{1},\dots,Y_{i-1}),$ (11)
where MAD is the median absolute deviation. MAD is defined as $\mathrm{MAD}=\operatorname{median}(|Y_{i}-\operatorname{median}(Y)|)$. These are robust to outliers and skewness. If $s_{i-1}$ is zero or not finite, we set $s_{i-1}=1$ to avoid degeneracy.
For the new patient, we form a standardized residual:
$r_{i}=\frac{Y_{i}-m_{i-1}}{s_{i-1}}.$ (12)
This means that the patient residual $r_{i}$ is the observed value ($Y_{i}$) minus the median observed so far ($s_{i-1}$ divided by the MAD ($s_{i-1}$). Note how this uses information for patients before $i$, keeping the martingale.
We then squash this standardized value into the interval (1,1)(-1,1) using
$g_{i}=\frac{r_{i}}{1+|r_{i}|}.$
This is a simple monotone transformation: for moderate values $g_{i}\approx r_{i}$, while very large positive or negative residuals are shrunk toward +1+1 or 1-1. The only purpose of this step is to prevent a single extreme observation from forcing an almost all-in bet.
Next, we ramp up the betting strength over time. Let
$c_{i}=\min\left\{1,\max\left(0,\frac{i-\text{burn-in}}{\text{ramp}}\right)\right\},$ (13)
where burn-in is the number of initial patients during which we essentially do not bet, and ramp controls how quickly we move from very cautious betting to our maximum aggressiveness. Those concepts are exactly like the binary approach. Finally, we cap the maximum betting strength at $c_{\max}\in(0,0.5]$ to avoid pathological bets.
The betting fraction $\lambda_{i}$ is then
$\lambda_{i}=0.5+c_{i}c_{\max}g_{i}.$ (14)
By construction, $\lambda_{i}\in(0,1)$ and is predictable: it depends only on past outcomes and the new $Y_{i}$, not on $T_{i}$.
Intuitively:
  • If $Y_{i}$ is close to the historical median, $g_{i}\approx 0$ and $\lambda_{i}\approx 0.5$: we essentially do not bet.
  • If $Y_{i}$ is unusually low (for example, very high ventilator-free days in a trial where we expect intervention to improve outcomes), then $g_{i}<0$ or $g_{i}>0$ depending on how we code the outcome and the expected direction of benefit. The sign of $g_{i}$ determines which arm we lean towards.
  • Early in the trial, $c_{i}$ is small, so even unusual observations lead to mild bets. As data accumulates, $c_{i}$ approaches 1 and the bets become more confident.
4.3 Wealth update
Treatment is still randomized with probability $p$ of intervention. After we choose $\lambda_{i}$, the wealth updates exactly as in the binary approach:
$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=1\ (\text{intervention})\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=0\ (\text{control})\end{cases}$ (15)
with $W_{0}=1$.
The only difference from the binary case is how we choose $\lambda_{i}$, as we saw. For binary outcomes, $\lambda_{i}$ depends on the event indicator and past event rates by arm. For continuous outcomes, $\lambda_{i}$ depends on how extreme $Y_{i}$ is relative to past outcomes.
4.4 Worked intuition
Imagine a trial where the outcome is ventilator-free days, and higher is better. Suppose that after 100 patients, the median and MAD of $Y$ are roughly stable. Patient 101 has an unusually high number of ventilator-free days compared with this distribution. The standardized residual $r_{101}$ is positive and large, so $g_{101}\approx 0.8$ and, after burn-in, $c_{101}\approx 1$. With $c_{\max}=0.6$, we might get $\lambda_{101}\approx 0.5+0.6\times 0.8\approx 0.98$: we strongly bet that this patient was in the intervention arm. If they indeed were, wealth increases by roughly a factor of $0.98/0.5\approx 2$ for this one patient. If not, wealth shrinks by $(1-0.98)/0.5\approx 0.04$.
Under the null, high values like this are just as likely in control as in intervention: we win and lose in balance, and wealth does not grow on average. Under a true benefit, such favorable outliers cluster in the intervention arm, and the bets pay off more often than not.
4.5 Validity
The key point is that validity does not depend on the choice of median, MAD, or the specific transformation $g_{i}$. It depends only on the fact that:
  1. 1.
    $\lambda_{i}$ is chosen before observing $T_{i}$ and depends only on past data and $Y_{i}$;
  2. 2.
    under the null, $T_{i}$ is independent of $Y_{i}$ with $\mathbb{P}(T_{i}=1)=p$.
Theorem 2.
Under the null hypothesis of no treatment effect, the e-process wealth process $(W_{i})$ is a test martingale: for all $i$,
$\mathbb{E}[W_{i}\mid\mathcal{F}_{i-1}]\leq W_{i-1},$
where $\mathcal{F}_{i-1}$ is the sigma-field generated by all observations up to step i1i-1.
Proof.
Condition on $\mathcal{F}_{i-1}$ and $Y_{i}$. The bet $\lambda_{i}$ is now fixed. Under the null, $T_{i}$ is independent of $Y_{i}$ and
$\mathbb{P}(T_{i}=1\mid\mathcal{F}_{i-1},Y_{i})=p,\quad\mathbb{P}(T_{i}=0\mid\mathcal{F}_{i-1},Y_{i})=1-p.$
The conditional expectation of the wealth multiplier is:
$\mathbb{E}\!\left[\frac{W_{i}}{W_{i-1}}\,\middle|\,\mathcal{F}_{i-1},Y_{i}\right]$ $=p\cdot\frac{\lambda_{i}}{p}+(1-p)\cdot\frac{1-\lambda_{i}}{1-p}$ (16)
$=\lambda_{i}+(1-\lambda_{i})$ (17)
$=1.$ (18)
Thus $\mathbb{E}[W_{i}\mid\mathcal{F}_{i-1},Y_{i}]=W_{i-1}$, and taking expectations over $Y_{i}$ yields $\mathbb{E}[W_{i}\mid\mathcal{F}_{i-1}]=W_{i-1}$. This shows that $(W_{i})$ is a martingale with unit expectation under the null. ∎
As in the binary case, Ville’s inequality implies that for any stopping time $\tau$,
$\mathbb{P}(W_{\tau}\geq 1/\alpha)\leq\alpha,$
so rejecting the null when $W_{\tau}\geq 1/\alpha$ controls Type I error at level $\alpha$, regardless of the stopping rule.
4.6 Simulation overview
We evaluated e-RTC using the same design philosophy as for binary approach. For a given standardized effect size (Cohen’s $d$) and target power, we first computed the fixed-sample size required for a two-sample $t$-test at $\alpha=0.05$. We then simulated trials with that sample size, assigning patients 1:1 to intervention or control, with outcomes drawn from normal distributions of equal variance and means differing by $d$ under the alternative.
The test was run sequentially with a burn-in period and ramp (for example, burn-in = 20 patients, ramp = 50 patients, $c_{\max}=0.6$). Under the null (no mean difference between arms), Type I error was close to or below the nominal level. Under the alternative, the implementation for continuous endpoints rejected the null with moderate power and typically crossed the threshold at an intermediate sample size (Table 2). Again, we believe that this may not be a replacement for the fixed-sample $t$-test but a conservative, anytime-valid monitoring tool that can trigger early stopping when effects are larger or clearer than anticipated.
Table 2: Operating Characteristics for Continuous Outcomes
Cohen’s $d$ Target Power $n$ Type I Error e-process Power Median Crossing
0.20 80% 788 0.042 14.1% 66
0.40 80% 200 0.038 33.6% 66
0.60 80% 90 0.037 55.1% 63
0.20 90% 1054 0.043 14.1% 67
0.40 90% 266 0.043 34.8% 66
0.60 90% 120 0.045 58.5% 65
A visual representation can be found in figure below.
Figure 3: Trajectories of the continuous randomization e-process for a trial designed to detect a standardized mean difference of $d=0.40$ with 80% power. Left: trajectories under the null hypothesis ($d=0$), where wealth wanders near or below 1 and rarely approaches the rejection threshold ($1/\alpha=20$). Right: trajectories under the alternative hypothesis ($d=0.40$), where wealth grows systematically and many paths cross the rejection threshold before the planned sample size is reached.
Figure 3 illustrates 30 simulated trajectories of the process for continuous endpoints under a design targeting a standardized effect size of $d=0.40$ with 80% power. Under the null hypothesis (left panel), wealth fluctuates around 1 and gradually drifts downward as repeated small bets accumulate against noise. No trajectory crosses the rejection threshold of $1/\alpha=20$. Under the corresponding alternative (right panel), outcomes begin to separate between arms, the bets become systematically correct, and the wealth grows. Some trajectories cross the rejection threshold before the planned sample size, demonstrating the potential for early stopping.
5 Time-to-Event Outcomes
Clinical trials often use time-to-event endpoints (e.g., overall survival), usually analyzed via the Log-Rank test or Cox proportional hazards models. These traditional methods require assumptions about proportional hazards or require waiting for a specific number of events. One can extend the randomization e-process to survival data, constructing a sequential Log-Rank test that updates wealth at every observed event.
pmlr-v146-grunwald21a developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio. We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model. We call this e-survival.
5.1 Setup and Martingale Construction
Let $N$ patients be randomized to treatment ($T=1$) or control ($T=0$). We observe outcomes over time. The "time" scale here is the distinct ordering of events. Let $t_{1}<t_{2}<\dots<t_{k}$ denote the times at which events occur.
At any event time $t_{j}$, we define the risk set $\mathcal{R}_{j}$ as the set of patients who have not yet had an event and have not been censored. Let $Y_{1}(t_{j})$ and $Y_{0}(t_{j})$ be the number of patients at risk in the treatment and control arms, respectively.
Under the null hypothesis of no treatment effect, the probability that the event at time $t_{j}$ comes from the treatment arm, conditional on a failure occurring within $\mathcal{R}_{j}$, is simply the proportion of treated patients at risk:
$p_{j}=\frac{Y_{1}(t_{j})}{Y_{1}(t_{j})+Y_{0}(t_{j})}.$ (19)
Let $X_{j}$ be the indicator that the event at $t_{j}$ is a treated patient ($X_{j}=1$ if treated, 0 if control). Under the null, $X_{j}$ is a Bernoulli trial with probability $p_{j}$. We construct the martingale increment (score) as:
$U_{j}=X_{j}-p_{j}.$ (20)
Note that $\mathbb{E}[U_{j}|\mathcal{R}_{j}]=0$.
5.2 Betting Strategy
We wager on the sign of $U_{j}$. If the treatment is beneficial, events will occur more slowly in the treatment arm than expected under the null. Thus, the observed number of treatment events will be lower than the expected number, leading to a negative trend in the cumulative sum of $U_{j}$.
We define the cumulative Log-Rank score at step j1j-1 as $Z_{j-1}=\sum_{k=1}^{j-1}U_{k}$. Our betting strategy targets this trend:
$\lambda_{j}=\text{sign}(Z_{j-1})\cdot c_{j}\cdot\lambda_{\max},$ (21)
where $c_{j}\in[0,1]$ is a ramping function similar to previous sections, and $\lambda_{\max}<1$ is a cap on betting aggressiveness.
The wealth update at event $j$ is:
$W_{j}=W_{j-1}\times(1+\lambda_{j}U_{j}).$ (22)
Because $\mathbb{E}[U_{j}]=0$, the expected multiplicative factor is 1 under the null. Thus, $(W_{j})$ is a test martingale.
Note that unlike the binary and continuous approaches, this betting strategy uses only the sign of the cumulative score $Z_{j-1}$, not its magnitude. Once evidence favors one direction, the bet size is fixed at $\lambda_{\max}$ regardless of how strong the accumulated evidence is. This mirrors the Kelly criterion in betting theory: the optimal wager size depends on the expected edge, and $\lambda_{\max}=0.25$ is calibrated for moderate effects (HR $\approx 0.80$(kelly1956). The connection between the wager, $\lambda$, and Kelly’s ideas on fraction of betting needs to be further explored. In brief, it makes sense that wager should be higher when prospects of winning are more favorable. For time-to-event, however, adapting response to events may take a long time, and a fixed wager may be preferable.
The parameters burn-in, ramp, and $\lambda_{\max}$ are set arbitrarily (burn-in =30=30, ramp =50=50, and $\lambda_{\max}=0.25$ in simulations). Different choices will yield different operating characteristics. The validity of the test does not depend on these choices—only efficiency does.
5.3 Handling Staggered Entry
In clinical practice, patients are recruited over time (staggered entry), whereas this simplified simulation generates survival times simultaneously. However, the Log-Rank test and this betting strategy rely solely on the rank ordering of events based on “time on study.”
We verified the validity of this simplification by simulating two scenarios with $N=631$ and a true Hazard Ratio of 0.80: (1) simultaneous entry, and (2) staggered entry where patients were recruited uniformly over 12 months, and analysis was performed on calculated study time ($T_{\text{event}}-T_{\text{entry}}$). The resulting distributions of final e-values were similar (Simultaneous Median $E\approx 25.1$, Staggered Median $E\approx 23.6$; Power $\approx 54\%$ and $53\%$ respectively). This confirms that the sequential e-process remains valid for real-world staggered designs provided the analysis utilizes time-since-randomization.
5.4 Simulation Results
We simulated survival trials comparing exponential survival times with a Hazard Ratio (HR) of 0.80. Targeted power was 80% using a standard Log-Rank design, which requires approximately 631 events. A total of $N=631$ patients (assuming no censoring) were used for simulations. Results are shown in Table 3.
Table 3: Operating Characteristics (N=631)
True HR Target HR Type I Error Power Median Events to Stop
1.00 (Null) 0.80 0.039
0.80 (Alt) 0.80 62.8% 329 (52% of N)
Under the alternative (HR=0.80), the e-survival process achieved 62.8% power to reject the null, with a median stopping time of 329 events. Examples are shown in Figure 4.
Figure 4: Trajectories of the e-Survival process for a trial designed to detect a Hazard Ratio of 0.80 with 80% power ($N=631$). Left: trajectories under the null hypothesis (HR = 1.00), where wealth fluctuates randomly. Right: trajectories under the alternative hypothesis (HR = 0.80), where wealth grows systematically. The red dashed line represents the rejection threshold ($1/\alpha=20$).
6 Discussion
The e-Rt is an nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a). The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses. It can be extended to continuous endpoints and time-to-event analyses.
6.1 Operating characteristics
Simulations demonstrate that the method properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
Both the continuous (e-RTC) and the survival (e-survival) implementations had similar characteristics: Power was traded for continuous Type 1 error control. All simulations showed that Type 1 error was exceptionally well controlled but power was nominally low.
These results should be interpreted carefully. From a clinical trial perspective, it is uncertain whether those methods can replace traditional frequentist or Bayesian paradigms, but they may provide a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule. If the e-process crosses its threshold, one could consider stopping the trial, for example. If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
6.2 What is the null hypothesis being tested?
This approach tests whether treatment assignment can be predicted from outcomes—equivalently, whether outcomes are exchangeable between arms. Under the null, $Y_{i}\perp T_{i}$ at each observation: knowing the outcome provides no information about which arm the patient belongs to. This is neither Fisher’s sharp null (every individual has exactly zero treatment effect) nor the weak null of equal population means.
A useful analogy is a casino. Under the null, the house is fair: no betting strategy can systematically grow wealth. The e-value quantifies accumulated evidence that the game is beatable. Rejecting the null means we have found a profitable strategy—outcomes predict assignments better than chance. The user is also referenced to the pioneer lessons by ramdas2021game.
This framing clarifies both the method’s strength and its limitation. The strength is generality: any departure from exchangeability—constant effects, heterogeneous effects, or even randomization failures—makes outcomes informative and wealth grows. The limitation is that the test detects only departures that a cumulative backward-looking betting strategy can exploit. In practical terms, the null hypothesis is best understood as: “there exists no stable, learnable relationship between outcomes and treatment assignment that persists long enough for an adaptive strategy to exploit.” This is appropriate for most trials where effects are expected to be consistent over enrollment, but investigators should be aware that non-stationary effects represent a blind spot.
6.3 Relationship to existing work
The betting framework for hypothesis testing was developed by shafer2021. E-values and e-processes have been extensively studied (vovk2021; ramdas2022; ramdas2025). pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which applies the betting framework directly to randomized experiments: an analyst sequentially bets on treatment assignments based on observed outcomes, with wealth forming a test martingale under the null. The binary approach implements this framework with a specific adaptive betting strategy tied to outcome values. Betting approaches have been established for estimating means of bounded random variables (waudbysmith2023estimating). The continuous extension adapts these principles to the two-sample randomization setting using a standardization strategy.
koning2025 develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations. pmlr-v146-grunwald21a developed the ‘Safe Log-rank Test’ based on evaluating likelihood ratios with specific priors on the hazard ratio to ensure growth rate optimality. In contrast, the survival approach constructs a linear test martingale directly from the score function using an adaptive betting strategy. Rather than relying on likelihood integration or specific priors, the survival approach process uses a heuristic ‘plug-in’ estimate of the effect direction, modulated by a ramping function. This offers a computationally simple, algorithmic alternative that derives validity strictly from the randomization probabilities within the risk set, without requiring the full apparatus of partial likelihood theory.
The binary e-RT shares the same martingale foundation as i-bet but differs in key respects: it operates prospectively as patients enroll rather than retrospectively on completed data; it requires no covariates or working models; and it uses betting fractions that adapt continuously to running outcome estimates rather than fixed-magnitude wagers guided by covariate-based predictions. This yields a simpler method that may be suitable for real-time trial monitoring.
It is possible that some of the concepts here were discussed by other authors in different contexts that were not immediately available for this author. Reader is encouraged to reach out if that is the case, and the author will happily adjust accordingly.
6.4 Limitations
Several limitations should be noted. This is an experimental method under development. Simulations are not exhaustive and were done for binary, continuous endpoints, and time to event. It is uncertain how this method would behave in more complex models, including competing risk models.
The methods proposed test only whether there are differences between arms. They do not directly estimate treatment effects or provide confidence intervals. The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates. Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power. Fourth, our simulations used specific betting strategies; other choices may yield different operating characteristics. Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power. Finally, it is unclear how this method will behave in situations where heterogeneity in treatment effects exist or there are temporal instabilities in effect size. The method could be extended to bet according to relative effect size approaches, such as the odds ratio. This is also under development.
6.5 Future Directions
Several extensions merit exploration. First, the current betting strategy uses a cumulative estimate $\hat{\delta}$ that weights all historical observations equally. This makes the method vulnerable to time-varying effects: if treatment benefit reverses to harm mid-trial, the strategy continues betting on stale information and wealth erodes despite continuous violation of exchangeability. Adaptive weighting schemes—such as exponential decay, rolling windows, or hybrid estimators blending long-term and recent signals—could improve robustness to non-stationary effects. Second, the betting intensity could itself adapt to recent performance: increasing $\lambda$ during sustained wealth growth (exploiting a confirmed edge) and dampening it following drawdowns (protecting against regime change). This mirrors Kelly criterion extensions that incorporate drawdown constraints. These refinements trade power under stable effects against robustness to drift, and the optimal balance likely depends on the clinical context. This is under development.
6.6 Conclusion
E-processes provide conceptually valid anytime-valid sequential inference for randomized trials using only the guarantee of randomization. Its validity is unconditional on the data-generating process, making it a potentially useful tool for trial monitoring. While it trades power for this robustness, it offers a valuable complement to traditional analysis methods.
7 Disclaimers and Version Control
7.1 Disclaimer
This is an experimental method under development. Application to real patients should only be considered under surveillance from an experienced statistician and remain strongly discouraged at this point by the author. The author is not responsible for consequences of use of this method.
7.2 LLM use statement
Large language models were extensively used in this work. The author had the idea that perhaps the e-value and e-process machinery could be used to bet against randomization which would result in a continuous trial monitoring tool. They uploaded the references in this manuscript to Gemini 3.0 Pro for brainstorming, which quickly resulted in a preliminary version. This was refined, tested, and debugged using Claude 4.5 Opus and ChatGPT 5.1 Pro. Gemini 3.0 Pro aided with coding for survival approach. Claude 4.5 Opus provided final spellchecking and organization of this draft, which remains suboptimal despite the efforts of several LLMs.
7.3 Acknowledgments
The author is thankful to Aaditya Ramdas for their thoughtful comments on the first version and for pointing out previous literature to the author.
7.4 Version Control
  1. 1.
    First Version (Dec 04, 2025)
  2. 2.
    Second Version (Dec 07, 2025): Minor text adjustments; removed claim on sharp null.
  3. 3.
    Third Version (Dec 11, 2025): Added continuous and survival endpoints; text adjustments.
  4. 4.
    Forth Version (This version): Correction on signal direction on -RTC. Text adjustments
Appendix A R Code
# e-RT Simulations: Type I Error and Power (Unified)
library(tidyverse)
# Core e-RT function
compute_eRT <- function(treatment, outcome, p = 0.5, burn_in = 50, ramp = 100) {
n <- length(treatment)
if (is.factor(treatment)) treatment <- as.numeric(treatment) - 1
if (is.factor(outcome)) outcome <- as.numeric(outcome) - 1
wealth <- numeric(n)
wealth[1] <- 1
for (i in 2:n) {
trt_prev <- treatment[1:(i-1)]
out_prev <- outcome[1:(i-1)]
rate_trt <- mean(out_prev[trt_prev == 1])
rate_ctrl <- mean(out_prev[trt_prev == 0])
if (is.nan(rate_trt)) rate_trt <- 0.5
if (is.nan(rate_ctrl)) rate_ctrl <- 0.5
delta_hat <- rate_trt - rate_ctrl
c_i <- max(0, min(1, (i - burn_in) / ramp))
if (outcome[i] == 1) {
lambda <- 0.5 + 0.5 * c_i * delta_hat
} else {
lambda <- 0.5 - 0.5 * c_i * delta_hat
lambda <- max(0.001, min(0.999, lambda))
if (treatment[i] == 1) {
multiplier <- lambda / p
} else {
multiplier <- (1 - lambda) / (1 - p)
wealth[i] <- wealth[i-1] * multiplier
return(wealth)
# Simulate a single trial
simulate_trial <- function(n, p_trt = 0.5, rate_trt, rate_ctrl) {
treatment <- rbinom(n, 1, p_trt)
outcome <- numeric(n)
outcome[treatment == 1] <- rbinom(sum(treatment == 1), 1, rate_trt)
outcome[treatment == 0] <- rbinom(sum(treatment == 0), 1, rate_ctrl)
return(data.frame(treatment = treatment, outcome = outcome))
# Unified simulation function
simulate_eRT <- function(n_sims = 5000,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100) {
# Determine hypothesized ARR for sample size calculation
if (is.null(hypothesized_ARR) && is.null(p_trt)) {
stop("Mustspecifyeitherp_trtorhypothesized_ARR")
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
# Calculate sample size based on hypothesized effect
$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
power = target_power, sig.level = alpha)
n_per_arm <- ceiling(ss\$n)
n_patients <- 2 * n_per_arm
# Determine true p_trt (null if not specified)
if (is.null(p_trt)) {
p_trt <- p_ctrl # Null is true
true_ARR <- 0
sim_type <- "TypeIError"
} else {
true_ARR <- p_ctrl - p_trt
sim_type <- "Power"
[…]
p_ctrl, hypothesized_ARR, target_power * 100))
[…]
p_ctrl, p_trt, true_ARR))
[…]
rejections <- 0
first_crossing <- numeric(n_sims)
final_evalues <- numeric(n_sims)
pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
for (sim in 1:n_sims) {
trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
final_evalues[sim] <- wealth[n_patients]
crossing <- which(wealth >= 1/alpha)
if (length(crossing) > 0) {
rejections <- rejections + 1
first_crossing[sim] <- crossing[1]
} else {
first_crossing[sim] <- NA
setTxtProgressBar(pb, sim)
close(pb)
rejection_rate <- rejections / n_sims
se <- sqrt(rejection_rate * (1 - rejection_rate) / n_sims)
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
if (sim_type == "Power") {
[…]
} else {
[…]
return(list(
sim_type = sim_type,
rejection_rate = rejection_rate,
se = se,
n_per_arm = n_per_arm,
n_patients = n_patients,
p_ctrl = p_ctrl,
p_trt = p_trt,
hypothesized_ARR = hypothesized_ARR,
true_ARR = true_ARR,
target_power = target_power,
alpha = alpha,
median_crossing = median_stop,
first_crossing = first_crossing,
final_evalues = final_evalues
# Plot trajectories
plot_trajectories <- function(n_trials = 30,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100,
title = NULL) {
if (is.null(hypothesized_ARR) && is.null(p_trt)) {
stop("Mustspecifyeitherp_trtorhypothesized_ARR")
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
if (is.null(p_trt)) {
p_trt <- p_ctrl # Null
# Calculate sample size
$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
power = target_power, sig.level = alpha)
n_patients <- 2 * ceiling(ss\$n)
true_ARR <- p_ctrl - p_trt
if (is.null(title)) {
if (true_ARR == 0) {
title <- sprintf("e-RTUnderNull(n=%d,designedfor%.0f%%ARR)",
n_patients, hypothesized_ARR * 100)
} else {
title <- sprintf("e-RTUnderAlternative(n=%d,%.0f%%power)",
n_patients, target_power * 100)
trajectories <- list()
for (i in 1:n_trials) {
trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
trajectories[[i]] <- data.frame(
patient = 1:n_patients,
wealth = wealth,
trial = i
df <- bind_rows(trajectories)
p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
geom_line(alpha = 0.4, color = "steelblue") +
geom_hline(yintercept = 1/alpha, linetype = "dashed", color = "red") +
geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
scale_y_log10() +
labs(
title = title,
subtitle = sprintf("Control=%.0f%%,Treatment=%.0f%%,TrueARR=%.0f%%",
p_ctrl * 100, p_trt * 100, true_ARR * 100),
x = "Patient",
y = "e-value(logscale)"
annotate("text", x = n_patients * 0.95, y = 1/alpha * 1.5,
label = sprintf("1/alpha=%.0f", 1/alpha), color = "red", hjust = 1) +
theme_minimal() +
theme(
plot.title = element_text(face = "bold"),
panel.grid.minor = element_blank()
return(p)
# Run all simulations
run_all <- function(n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05,
seed = 42) {
set.seed(seed)
results <- list()
# Create grid of scenarios
scenarios <- expand.grid(
hypothesized_ARR = ARRs,
target_power = target_powers,
stringsAsFactors = FALSE
# For each scenario, run Type I error and Power
all_results <- data.frame()
for (i in 1:nrow(scenarios)) {
hyp_ARR <- scenarios\$hypothesized_ARR[i]
tgt_pow <- scenarios\$target_power[i]
[…]
hyp_ARR * 100, tgt_pow * 100))
[…]
# Type I error (null true, same sample size)
[…]
t1 <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = NULL, # Null is true
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
alpha = alpha
# Power (alternative true)
[…]
pow <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = p_ctrl - hyp_ARR, # True effect = hypothesized
target_power = tgt_pow,
alpha = alpha
all_results <- rbind(all_results, data.frame(
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
n_patients = t1\$n_patients,
type1_error = t1\$rejection_rate,
type1_se = t1\$se,
eRT_power = pow\$rejection_rate,
power_se = pow\$se,
median_crossing = pow\$median_crossing
results\$summary <- all_results
# Print summary
[…]
print(all_results %>%
mutate(
hypothesized_ARR = sprintf("%.0f%%", hypothesized_ARR * 100),
target_power = sprintf("%.0f%%", target_power * 100),
type1_error = sprintf("%.3f", type1_error),
eRT_power = sprintf("%.1f%%", eRT_power * 100)
select(hypothesized_ARR, target_power, n_patients,
type1_error, eRT_power, median_crossing))
return(results)
# Run
if (interactive()) {
results <- run_all(
n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05
# Trajectory plots - Alternative
p1 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.80)
ggsave("traj_alt_10pct_80pow.pdf", p1, width = 8, height = 5)
p2 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.90)
ggsave("traj_alt_10pct_90pow.pdf", p2, width = 8, height = 5)
# Trajectory plots - Null
p3 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.80)
ggsave("traj_null_10pct_80pow.pdf", p3, width = 8, height = 5)
p4 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.90)
ggsave("traj_null_10pct_90pow.pdf", p4, width = 8, height = 5)
# e-RTC (Randomization E-Process for Continuous Outcomes)
# Type I Error and Power Simulations (Unified)
setwd("")
suppressPackageStartupMessages({
library(tidyverse)
# Core e-RTC function
# treatment: 0/1 vector (0 = control, 1 = intervention)
# outcome : continuous outcome (numeric)
# p : allocation prob for treatment (default 0.5)
# burn_in : number of patients before bets start ramping up
# ramp : additional patients over which betting strength ramps to max
# c_max : maximum betting strength (0 < c_max <= 0.5 recommended)
compute_eRTC <- function(treatment, outcome,
p = 0.5,
burn_in = 20,
ramp = 50,
c_max = 0.6) {
n <- length(treatment)
if (is.factor(treatment)) treatment <- as.numeric(treatment) - 1
wealth <- numeric(n)
wealth[1] <- 1
for (i in 2:n) {
# Past outcomes
out_prev <- outcome[1:(i-1)]
# Require enough past data before betting
if ((i - 1) < burn_in) {
wealth[i] <- wealth[i-1]
next
# Robust center and scale: median + MAD
med_prev <- median(out_prev, na.rm = TRUE)
mad_prev <- mad(out_prev, center = med_prev, constant = 1, na.rm = TRUE)
if (!is.finite(mad_prev) $|$ mad_prev <= 0) mad_prev <- 1
# Current outcome
y_i <- outcome[i]
# Standardized residual, squashed to (-1, 1)
r_i <- y_i - med_prev
s_i <- r_i / mad_prev
g_i <- s_i / (1 + abs(s_i)) # in (-1,1), robust to outliers
# Learn direction from past data (NEW)
trt_prev <- treatment[1:(i-1)]
mean_trt <- mean(out_prev[trt_prev == 1])
mean_ctrl <- mean(out_prev[trt_prev == 0])
# Handle edge cases early in trial
if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) {
direction <- 0
} else {
direction <- sign(mean_trt - mean_ctrl)
# Ramping betting strength
ramp_frac <- max(0, min(1, (i - burn_in) / ramp))
c_i <- ramp_frac * c_max
lambda <- 0.5 + c_i * g_i * direction
lambda <- max(0.001, min(0.999, lambda))
# Wealth update based on actual assignment
if (treatment[i] == 1) {
multiplier <- lambda / p
} else {
multiplier <- (1 - lambda) / (1 - p)
wealth[i] <- wealth[i-1] * multiplier
return(wealth)
# Simulate a single trial with continuous outcomes
# n : total number of patients
# p_trt : probability of being assigned to treatment (default 0.5)
# mu_ctrl : mean in control arm
# mu_trt : mean in treatment arm
# sd : common standard deviation (default 1)
simulate_trial_continuous <- function(n,
p_trt = 0.5,
mu_ctrl = 0,
mu_trt = 0,
sd = 1) {
treatment <- rbinom(n, 1, p_trt)
outcome <- numeric(n)
outcome[treatment == 0] <- rnorm(sum(treatment == 0), mean = mu_ctrl, sd = sd)
outcome[treatment == 1] <- rnorm(sum(treatment == 1), mean = mu_trt, sd = sd)
return(data.frame(treatment = treatment, outcome = outcome))
# Unified simulation function for e-RTC
# n_sims : number of simulated trials
# mu_ctrl : control mean (default 0)
# mu_trt : treatment mean (NULL => null, mu_trt = mu_ctrl)
# hypothesized_d: Cohen’s d used for sample size calc (delta/sd)
# target_power : design power (0.80, 0.90)
# alpha : significance level
# burn_in, ramp, c_max: e-RTC tuning parameters
simulate_eRTC <- function(n_sims = 5000,
mu_ctrl = 0,
mu_trt = NULL,
hypothesized_d,
target_power = 0.80,
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6) {
if (missing(hypothesized_d)) {
stop("Mustspecifyhypothesized_d(Cohen’sdfordesign).")
# Treat SD = 1 for design (Cohen’s d = delta / sd)
sd_true <- 1
delta_design <- hypothesized_d * sd_true
# Sample size via t-test (two-sample, equal sizes)
$s$ <- power.t.test(delta = delta_design,
sd = sd_true,
power = target_power,
sig.level = alpha,
type = "two.sample",
alternative = "two.sided")
n_per_arm <- ceiling(ss\$n)
n_patients <- 2 * n_per_arm
# Determine true mu_trt (null if not specified)
if (is.null(mu_trt)) {
mu_trt <- mu_ctrl # Null is true
true_d <- 0
sim_type <- "TypeIError"
} else {
true_d <- (mu_trt - mu_ctrl) / sd_true
sim_type <- "Power"
[…]
mu_ctrl, hypothesized_d, target_power * 100))
[…]
mu_ctrl, mu_trt, true_d))
[…]
rejections <- 0
first_crossing <- numeric(n_sims)
final_evalues <- numeric(n_sims)
pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
for (sim in 1:n_sims) {
trial <- simulate_trial_continuous(n_patients,
mu_ctrl = mu_ctrl,
mu_trt = mu_trt,
sd = sd_true)
wealth <- compute_eRTC(trial\$treatment,
trial\$outcome,
p = 0.5,
burn_in = burn_in,
ramp = ramp,
c_max = c_max)
final_evalues[sim] <- wealth[n_patients]
crossing <- which(wealth >= 1 / alpha)
if (length(crossing) > 0) {
rejections <- rejections + 1
first_crossing[sim] <- crossing[1]
} else {
first_crossing[sim] <- NA
setTxtProgressBar(pb, sim)
close(pb)
rejection_rate <- rejections / n_sims
se <- sqrt(rejection_rate * (1 - rejection_rate) / n_sims)
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
rejection_rate - 1.96 * se,
rejection_rate + 1.96 * se))
if (sim_type == "Power") {
[…]
} else {
[…]
return(list(
sim_type = sim_type,
rejection_rate = rejection_rate,
se = se,
n_per_arm = n_per_arm,
n_patients = n_patients,
mu_ctrl = mu_ctrl,
mu_trt = mu_trt,
hypothesized_d = hypothesized_d,
true_d = true_d,
target_power = target_power,
alpha = alpha,
median_crossing = median_stop,
first_crossing = first_crossing,
final_evalues = final_evalues
# Plot trajectories under null / alternative
# n_trials : number of trajectories to overlay
# mu_ctrl, mu_trt, hypothesized_d, target_power, alpha, burn_in, ramp, c_max
plot_trajectories_eRTC <- function(n_trials = 30,
mu_ctrl = 0,
mu_trt = NULL,
hypothesized_d,
target_power = 0.80,
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6,
title = NULL) {
if (missing(hypothesized_d)) {
stop("Mustspecifyhypothesized_d(Cohen’sd).")
sd_true <- 1
delta_design <- hypothesized_d * sd_true
$s$ <- power.t.test(delta = delta_design,
sd = sd_true,
power = target_power,
sig.level = alpha,
type = "two.sample",
alternative = "two.sided")
n_patients <- 2 * ceiling(ss\$n)
if (is.null(mu_trt)) {
mu_trt <- mu_ctrl # Null
true_d <- 0
} else {
true_d <- (mu_trt - mu_ctrl) / sd_true
if (is.null(title)) {
if (true_d == 0) {
title <- sprintf(
"e-RTCUnderNull(n=%d,designedford=%.2f)",
n_patients, hypothesized_d
} else {
title <- sprintf(
"e-RTCUnderAlternative(n=%d,trued=%.2f,%.0f%%powerdesign)",
n_patients, true_d, target_power * 100
trajectories <- vector("list", n_trials)
for (i in 1:n_trials) {
trial <- simulate_trial_continuous(n_patients,
mu_ctrl = mu_ctrl,
mu_trt = mu_trt,
sd = sd_true)
wealth <- compute_eRTC(trial\$treatment,
trial\$outcome,
p = 0.5,
burn_in = burn_in,
ramp = ramp,
c_max = c_max)
trajectories[[i]] <- data.frame(
patient = 1:n_patients,
wealth = wealth,
trial = i
df <- bind_rows(trajectories)
p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
geom_line(alpha = 0.4, color = "steelblue") +
geom_hline(yintercept = 1 / alpha, linetype = "dashed", color = "red") +
geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
scale_y_log10() +
labs(
title = title,
subtitle = sprintf("mu_ctrl=%.2f,mu_trt=%.2f,trued=%.2f",
mu_ctrl, mu_trt, true_d),
x = "Patient",
y = "e-value(logscale)"
annotate("text",
x = n_patients * 0.95,
y = (1 / alpha) * 1.5,
label = sprintf("1/alpha=%.0f", 1 / alpha),
color = "red", hjust = 1) +
theme_minimal() +
theme(
plot.title = element_text(face = "bold"),
panel.grid.minor = element_blank()
return(p)
# Run all scenarios (grid over d and power)
# ds : vector of Cohen’s d (e.g. c(0.2, 0.4, 0.6))
# target_powers: vector of target powers (e.g. c(0.80, 0.90))
run_all_eRTC <- function(n_sims = 5000,
mu_ctrl = 0,
ds = c(0.2, 0.4, 0.6),
target_powers = c(0.80, 0.90),
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6,
seed = 42) {
set.seed(seed)
results <- list()
scenarios <- expand.grid(
hypothesized_d = ds,
target_power = target_powers,
stringsAsFactors = FALSE
all_results <- data.frame()
for (i in 1:nrow(scenarios)) {
hyp_d <- scenarios\$hypothesized_d[i]
tgt_pow <- scenarios\$target_power[i]
[…]
hyp_d, tgt_pow * 100))
[…]
# Type I error
[…]
t1 <- simulate_eRTC(
n_sims = n_sims,
mu_ctrl = mu_ctrl,
mu_trt = NULL, # Null true
hypothesized_d = hyp_d,
target_power = tgt_pow,
alpha = alpha,
burn_in = burn_in,
ramp = ramp,
c_max = c_max
# Power
[…]
pow <- simulate_eRTC(
n_sims = n_sims,
mu_ctrl = mu_ctrl,
mu_trt = mu_ctrl + hyp_d, # sd = 1, so delta = d
hypothesized_d = hyp_d,
target_power = tgt_pow,
alpha = alpha,
burn_in = burn_in,
ramp = ramp,
c_max = c_max
all_results <- rbind(all_results, data.frame(
hypothesized_d = hyp_d,
target_power = tgt_pow,
n_patients = t1\$n_patients,
type1_error = t1\$rejection_rate,
type1_se = t1\$se,
eRTC_power = pow\$rejection_rate,
power_se = pow\$se,
median_crossing = pow\$median_crossing
results\$summary <- all_results
[…]
print(all_results %>%
mutate(
hypothesized_d = sprintf("%.2f", hypothesized_d),
target_power = sprintf("%.0f%%", target_power * 100),
type1_error = sprintf("%.3f", type1_error),
eRTC_power = sprintf("%.1f%%", eRTC_power * 100)
select(hypothesized_d, target_power, n_patients,
type1_error, eRTC_power, median_crossing))
return(results)
# Example run (interactive)
if (interactive()) {
# Grid of d and power
results_eRTC <- run_all_eRTC(
n_sims = 2000, # lower for quick tests; bump to 5000+ for final
mu_ctrl = 0,
ds = c(0.2, 0.4, 0.6),
target_powers = c(0.80, 0.90),
alpha = 0.05,
burn_in = 20,
ramp = 50,
c_max = 0.6
# Trajectory plots - Alternative examples
p_alt_04_80 <- plot_trajectories_eRTC(
mu_ctrl = 0,
mu_trt = 0.4, # true d = 0.4
hypothesized_d = 0.4,
target_power = 0.80,
alpha = 0.05
ggsave("traj_eRTC_alt_d0.4_80pow.pdf", p_alt_04_80, width = 8, height = 5)
# Trajectory plots - Null
p_null_04_80 <- plot_trajectories_eRTC(
mu_ctrl = 0,
mu_trt = NULL, # null
hypothesized_d = 0.4,
target_power = 0.80,
alpha = 0.05
ggsave("traj_eRTC_null_d0.4_80pow.pdf", p_null_04_80, width = 8, height = 5)
# e-Survival: Sequential Log-Rank Test (Unified)
library(tidyverse)
library(survival) # Used for standard verification if needed
# Core e-Survival function
# time: vector of event times
# status: vector of status (1=event, 0=censored)
# treatment: vector of assignments (1=trt, 0=ctrl)
# burn_in: events to observe before betting
# ramp: number of events to ramp up betting intensity
# lambda: maximum betting scalar (0 < lambda < 1)
compute_eSurvival <- function(time, status, treatment,
burn_in = 20,
ramp = 50,
lambda_max = 0.2) {
n <- length(time)
# 1. Create a dataframe and sort by time (events happen sequentially)
df <- data.frame(time, status, treatment) %>%
arrange(time)
# Vectors for processing
T_sorted <- df\$time
S_sorted <- df\$status
A_sorted <- df\$treatment
wealth <- numeric(n)
wealth[1] <- 1
# Running Log-Rank Score (Observed - Expected)
cumulative_Z <- 0
# Track active Risk Set manually for speed/simplicity
# Initially, everyone is at risk (assuming time 0 start)
risk_trt <- sum(treatment == 1)
risk_ctrl <- sum(treatment == 0)
# We need to process events one by one.
# Note: This simple loop assumes no tied times for clarity.
# (Ties are rare in continuous simulation).
for (i in 1:n) {
# --- A. Betting Step (Before observing outcome i) ---
# Determine bet direction based on history
if (i > burn_in) {
# Ramping factor
c_i <- max(0, min(1, (i - burn_in) / ramp))
# Direction:
# If cumulative_Z is negative, Treatment has FEWER events than expected (Benefit).
# We bet that this trend continues.
# If Observed (1) - Expected (p) is likely negative, we want a NEGATIVE bet
# to make the wealth multiplier (1 + bet * U) > 1.
# Simple strategy: proportional to the sign of the current trend
bet_direction <- sign(cumulative_Z)
if (bet_direction == 0) bet_direction <- 0
# The bet:
b_i <- c_i * lambda_max * bet_direction
} else {
b_i <- 0
# --- B. Observation Step ---
is_event <- S_sorted[i] == 1
tr_is_event <- (A_sorted[i] == 1)
# Probability that *if* an event happens, it is in Treatment arm
# p_null = (Risk Treatment) / (Total Risk)
total_risk <- risk_trt + risk_ctrl
if (total_risk > 0) {
p_null <- risk_trt / total_risk
} else {
p_null <- 0.5
# --- C. Update Wealth (Only on events, not censorings) ---
if (is_event) {
# U_i: The "score" contribution (Observed - Expected)
# If Treatment event: 1 - p
# If Control event: 0 - p
obs <- ifelse(tr_is_event, 1, 0)
U_i <- obs - p_null
# Martingale update
# Constraint: ensure (1 + b_i * U_i) is non-negative.
# Since |U_i| < 1 and lambda_max < 1, this is generally safe.
multiplier <- 1 + b_i * U_i
# Update history
cumulative_Z <- cumulative_Z + U_i
if (i > 1) {
wealth[i] <- wealth[i-1] * multiplier
} else {
wealth[i] <- multiplier
} else {
# Censored event: No bet made, wealth carries forward
if (i > 1) wealth[i] <- wealth[i-1]
# --- D. Update Risk Set for next step ---
# The person at index i is removed from risk set (event or censored)
if (tr_is_event) {
risk_trt <- max(0, risk_trt - 1)
} else {
risk_ctrl <- max(0, risk_ctrl - 1)
# Map wealth back to original unsorted indices is tricky in sequential plots,
# so we return the Sequential Wealth (wealth over event time).
return(data.frame(
event_num = 1:n,
time = T_sorted,
wealth = wealth
# Simulate a single survival trial (Weibull)
# HR < 1 implies Treatment is better (longer survival)
simulate_trial_survival <- function(n,
HR = 1,
shape = 1.2,
scale = 10,
cens_prop = 0.0) {
# 1. Treatments
treatment <- rbinom(n, 1, 0.5)
# 2. True Times (Weibull AFT)
# HR in Cox model corresponds roughly to scale change in Weibull
# Scale_trt = Scale_ctrl / (HR^(1/shape))
scale_trt <- scale / (HR^(1/shape))
U <- runif(n)
true_time <- numeric(n)
true_time[treatment == 0] <- scale * (-log(U[treatment==0]))^(1/shape)
true_time[treatment == 1] <- scale_trt * (-log(U[treatment==1]))^(1/shape)
# 3. Censoring
# Simple independent uniform censoring to achieve approx proportion
if (cens_prop > 0) {
C <- runif(n, 0, 2 * scale) # broad censoring distribution
# Adjust range to hit target prop roughly if needed, keeping simple here
time <- pmin(true_time, C)
status <- as.numeric(true_time <= C)
} else {
time <- true_time
status <- rep(1, n)
return(list(time = time, status = status, treatment = treatment))
# Unified simulation function
simulate_eSurvival <- function(n_sims = 1000,
n_patients = NULL, # Change to NULL to trigger auto-calc
HR_true = 1,
target_HR = 0.7,
target_power = 0.80, # Added for flexibility
alpha = 0.05,
burn_in = 30,
ramp = 50,
lambda_max = 0.25) {
# Determine simulation type
if (HR_true == 1) {
sim_type <- "TypeIError(Null)"
} else {
sim_type <- "Power(Alternative)"
# Calculate Sample Size if not provided
if (is.null(n_patients)) {
# Schoenfeld formula for events
z_alpha <- qnorm(1 - alpha/2)
z_beta <- qnorm(target_power)
# Total events needed
req_events <- ceiling( 4 * ((z_alpha + z_beta) / log(target_HR))^2 )
# Assuming no censoring, Patients = Events
n_patients <- req_events
[…]
target_power*100, target_HR, n_patients))
[…]
rejections <- 0
first_crossing <- numeric(n_sims)
pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
for (sim in 1:n_sims) {
trial <- simulate_trial_survival(n_patients, HR = HR_true)
res <- compute_eSurvival(trial\$time, trial\$status, trial\$treatment,
burn_in = burn_in, ramp = ramp, lambda_max = lambda_max)
# Check stopping
crossing <- which(res\$wealth >= 1/alpha)
if (length(crossing) > 0) {
rejections <- rejections + 1
first_crossing[sim] <- crossing[1]
} else {
first_crossing[sim] <- NA
setTxtProgressBar(pb, sim)
close(pb)
rate <- rejections / n_sims
se <- sqrt(rate * (1 - rate) / n_sims)
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
if (sim_type == "Power(Alternative)") {
[…]
return(list(rate = rate, median_stop = median_stop, n_patients = n_patients))
# Plot Trajectories
plot_trajectories_surv <- function(n_trials = 20,
n_patients = 400,
HR_true = 0.7,
alpha = 0.05,
title = "e-SurvivalTrajectories") {
trajectories <- list()
for (i in 1:n_trials) {
trial <- simulate_trial_survival(n_patients, HR = HR_true)
res <- compute_eSurvival(trial\$time, trial\$status, trial\$treatment)
res\$trial <- i
trajectories[[i]] <- res
df <- bind_rows(trajectories)
ggplot(df, aes(x = event_num, y = wealth, group = trial)) +
geom_line(alpha = 0.5, color = "darkgreen") +
geom_hline(yintercept = 1/alpha, linetype = "dashed", color = "red") +
scale_y_log10() +
labs(
title = title,
subtitle = sprintf("TrueHR=%.2f", HR_true),
x = "NumberofEvents",
y = "e-value(logscale)"
annotate("text", x = n_patients, y = 1/alpha,
label = sprintf("Threshold%.0f", 1/alpha), vjust = -0.5, color="red") +
theme_minimal()
# Run Scenarios
if (interactive()) {
# Define design parameters
target_hr <- 0.80
# 1. Check Type I Error (Null: HR = 1.0)
# Passing n_patients = NULL triggers the auto-calculation based on target_HR
simulate_eSurvival(
n_sims = 2000,
n_patients = NULL, # Auto-calculate N for 80% power
HR_true = 1.0, # Reality: Null
target_HR = target_hr,
alpha = 0.05
# 2. Check Power (Alternative: HR = 0.7)
simulate_eSurvival(
n_sims = 2000,
n_patients = NULL, # Auto-calculate N for 80% power
HR_true = 0.8, # Reality: Alternative works
target_HR = target_hr,
alpha = 0.05
# 3. Plot Trajectories
# We calculate N explicitly here just to pass it to the plot function
req_events <- ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(target_hr))^2)
p <- plot_trajectories_surv(
n_trials = 30,
n_patients = req_events,
HR_true = 0.8,
alpha = 0.05,
title = sprintf("e-Survival(TargetHR=%.2f,N=%d)", target_hr, req_events)
print(p)
set.seed(2)
p1<-plot_trajectories_surv(
n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
HR_true = 0.8)
ggsave("traj_survivaleRT_08.pdf",p1,
width = 8, height = 5)
p2<-plot_trajectories_surv(
n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
HR_true = 1)
p2
ggsave("traj_survivaleRT_null.pdf",p2,
width = 8, height = 5)
# APPENDIX: Verification of Staggered vs. Simultaneous Entry
# This script demonstrates that the "Simultaneous Entry" assumption used in the
# main simulation yields statistically identical results to a realistic
# "Staggered Entry" design, provided the analysis is based on "Time on Study".
verify_staggered_equivalence <- function(n_sims = 1000,
n_patients = 631, # Updated to match HR=0.8 scenario
HR_true = 0.8, # Updated to match HR=0.8 scenario
recruit_period = 12, # Recruitment over 12 months
alpha = 0.05) {
[…]
# Storage for final e-values
final_e_simultaneous <- numeric(n_sims)
final_e_staggered <- numeric(n_sims)
pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
for (i in 1:n_sims) {
# 1. Generate the underlying biology (Survival Times)
# -- Method A: Simultaneous (The Simple Way) --
trial_sim <- simulate_trial_survival(n_patients, HR = HR_true)
res_sim <- compute_eSurvival(trial_sim\$time, trial_sim\$status, trial_sim\$treatment)
final_e_simultaneous[i] <- tail(res_sim\$wealth, 1)
# -- Method B: Staggered (The "Real World" Way) --
# We generate fresh biology for the staggered arm to compare distributions
trial_stag <- simulate_trial_survival(n_patients, HR = HR_true)
# ... but we add "Recruitment Time" (Uniform over 12 months)
entry_time <- runif(n_patients, 0, recruit_period)
# The event happens at this Calendar Time:
calendar_event_time <- entry_time + trial_stag\$time
# The analyst calculates "Time on Study" (Analysis Time):
# Analysis Time = Calendar Event Time - Entry Time
calculated_study_time <- calendar_event_time - entry_time
# Run e-process on the calculated study time
# Note: compute_eSurvival handles the sorting of events internally
res_stag <- compute_eSurvival(calculated_study_time,
trial_stag\$status,
trial_stag\$treatment)
final_e_staggered[i] <- tail(res_stag\$wealth, 1)
setTxtProgressBar(pb, i)
close(pb)
# 2. Compare Results
df_res <- data.frame(
e_value = c(final_e_simultaneous, final_e_staggered),
Method = rep(c("Simultaneous(Simplified)", "Staggered(Realistic)"), each = n_sims)
# Summary Stats
print(df_res %>%
group_by(Method) %>%
summarise(
Mean_E = mean(e_value),
Median_E = median(e_value),
Power_Estimated = mean(e_value >= 1/alpha)
# 3. Visualization
p <- ggplot(df_res, aes(x = e_value, fill = Method)) +
geom_density(alpha = 0.5, color = NA) +
scale_x_log10() +
labs(
title = "EquivalenceofSimultaneousvs.StaggeredEntry",
subtitle = sprintf("Distributionoffinale-values(N=%d,HR=%.2f)", n_patients, HR_true),
x = "Finale-value",
y = "Density"
theme_minimal() +
theme(legend.position = "bottom")
return(p)
# Run the verification
if (interactive()) {
# Matching the N=631 / HR=0.8 scenario from the paper
p_verify <- verify_staggered_equivalence(
n_sims = 1000,
n_patients = 631,
HR_true = 0.8
print(p_verify)
# Optional: Save for supplement
ggsave("supp_staggered_equivalence.pdf", p_verify, width = 8, height = 5)